The sides of a triangle are consecutive multiples of 4. Its perimeter is 72. Find the length of its shortest side.
(a) 5 (b) 20 (c) 23 (d) 24 (e) 28

Found 2 solutions by subudear, richwmiller:
Answer by subudear(62)

answer is (b) 20
each side is 20, 24 ,28

Answer by richwmiller(17219)

4n+4*(n+1)+4*(n+2)=72
4*(n+n+1+n+2)=72
4*(3n+3)=72
12n+12=72
12n=60
n=5
4n=20